Pick out the solution from the values given in the bracket next to each equation. Show that the other values do not satisfy the equation.
n + 12                        ( 12, 8, 20, 0)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

n + 12 = 20
n = 8 is a solution for this equation because for n = 8,
n + 12 = 8 + 12
= 20
∴Equation satisfied
n = 12 is not a solution for this equation because for n = 12,
n + 12 = 12 + 12
= 24 and not 20
n = 20 is not a solution for this equation because for n = 20,
n + 12 = 20 + 12
= 32 and not 20
n = 0 is not a solution for this equation because for n = 0
n + 12 = 0 + 12
= 12 and not 20

Thanks